OpenStudy (anonymous):

I am having a horrible time with my math homework. Can someone try to explain this to me? (6,-9);9x-2y=3 I think I am suppose to find the slope and intercept.

OpenStudy (amistre64):

(x,y) so; x=6, and y=-9. The equation you have needs to be solved for y in order to find the slope and y-intercept. 9x -2y =3 -2y =-9x +3 y = (9/2)x - (3/2) Your yintercept is the value for "y" when "x=0": (-3/2) The slope of the line is whatever is sitting next to the "x"; which in this case is (9/2) Does (x=6, y=-9) make a true statement tho: 9(6) -2(-9) = 3? 54 + 18 = 3? False. So your given value (6,-9) are not a solution.

OpenStudy (amistre64):

The information you are given suggests that you are wanting to determine if (6,-9) is a solution to the equation (9x -2y =3). So just plug in (x=6, y=-9) and see if it is a true statement....

OpenStudy (anonymous):

Thanks